OpenStudy (anonymous):

Write an equation that represents the problem, then solve the problem. Seven-fourths decreased by a number is 2. Find the number 1/4 (7-4 x) = 2 right

OpenStudy (anonymous):

\(\large\frac{7}{4}-x=2\) 7-4x=8 4x=-1 x=-\(\large \frac{1}{4}\)

OpenStudy (mathstudent55):

Let \(x\) represent the number. \( \dfrac{7}{4} - x = 2\) Factoring out \(\dfrac{1}{4} \), you get: \( \dfrac{1}{4}( 7 - 4x ) = 2 \) Your equation is correct.